Grade 12th passMechanics

Starting a race , a 68.2 kg sprinter runs the first 7.04 m in 1.60 sec,starting from the rest and acceleration uniformly.
(A)what is the sprinter speed at the end of the 1.60s?
(B)what is the sprinter kinetic energy?
(C)what average power does the sprinter generate during 1.60 s interval?

Amama
7 Years agoGrade 12th pass

2 Answers

akshay
7 Years ago
(A)
  s=0.5 a t^2
 7.04=0.5 a 1.6^2
 a=5.5
 
v=at
 
v=5.5*1.6
.
v=8.8 m/s
 
 
(B)
    K.E=0.5mv^2
   K.E=0.5*68.2*(5.5*t)^2
   K.E=11346.775 * t^2 J
          at any time t
 
(C)
  P=m*a*s/t
 P=68.2*5.5*7.04/1.6
P=1650.44 J/s
Klare Montefalco
5 Years ago
An elevator with a 10 person capacity ( each person with an average mass of 68kg) raises passengers to a height of 152 m in 55.0s at constant speed . Find the power output of the motor assuming no frictional losses ?